Grade 12th passPhysical Chemistry

NH3 oxidation number.........??????
A) -1
B) +1
C) 0
D) -3

Harshit Singh
5 Years agoGrade 12th pass

1 Answer

Pawan Prajapati
5 Years ago
Answer: C) 0
Explanation:
NH3, Ammonia is a neutral compound as the individual oxidation numbers elements that make up the compound NH3 are Nitrogen (N) and Hydrogen (H) sum to zero.
We do not speak of the oxidation number of a molecule. We can speak of the oxidation numbers of the individual elements that make up the compound.
Here the oxidation number of Nirogen is -3 and that of Hydrogen is +1.
Now, NH3 = 1(-3) + 3(+1) = -3 + 3 = 0


Thanks